OpenStudy (anonymous):

Match the values of x with the solution for y if y plus two and five−twelfths equals x. Match Term Definition x equals one−half A) y equals negative four and one−fourth x equals three and one−third B) y equals negative three and one−sixth x equals negative three−fourths C) y equals eleven−twelfths x equals negative one and five−sixths D) y equals negative one and eleven−twelfths

OpenStudy (debbieg):

I can't be here long so please don't feel deserted if I am not answering... :) You have a "rule" that relates x and y:\[\Large x=y+2\frac{ 5 }{12 }\] So just plug in one of the variables (y is probably easiest for these) and solve for the other. E.g., for the first y given: \[\Large x=-4\frac{ 1 }{4 }+2\frac{ 5 }{12 }\] Evaluate that right hand side, and you'll have your x.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

y + 2 5/12 = x (change the mixed number to improper fraction) y + 29/12 = x x = 1/2 y + 29/12 = x y + 29/12 = 1/2 y = -29/12 + 1/2 (common denominator is 12) y = -29/12 + 6/12 y = - 23/12 = - 1 11/12 (D) you see how I did that ?

OpenStudy (anonymous):

now if x = 3 1/3 (change to improper fraction) x = 10/3 now sub 10/3 in for x y + 29/12 = x y + 29/12 = 10/3 y = -29/12 + 10/3 (common denominator is 12) y = -29/12 + 40/12 y = 11/12 (C)

OpenStudy (anonymous):

THANKS SO MUCH!

OpenStudy (anonymous):

can you do the other two ?

OpenStudy (anonymous):

tag me if you need help with them :)